evaluate ∫ {(x+1) (x – 2)}/sqrt(x) dx

evaluate
∫ {(x+1) (x – 2)}/sqrt(x) dx

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

I = \int \frac{(x+1)(x-2)}{\sqrt{x}}dx
I = \int \frac{x^2-x-2}{\sqrt{x}}dx
I = \int (x^{3/2}-\sqrt{x}+2x^{-1/2})dx
I = \int x^{3/2}dx-\int \sqrt{x}+\int 2x^{-1/2}dx
I = \frac{2x^{5/2}}{5}-\int \sqrt{x}+\int 2x^{-1/2}dx
I = \frac{2x^{5/2}}{5}-\frac{2x^{3/2}}{3}+\int 2x^{-1/2}dx
I = \frac{2x^{5/2}}{5}-\frac{2x^{3/2}}{3}+4\sqrt{x}+c